OpenStudy (anonymous):

Find the sum of the series when it converges. (if geometric, find the first term, the ratio and the sum)

OpenStudy (anonymous):

\[\sum_{\infty}^{n=0}2(\frac{ 2 }{ 3 })^{n}\]

OpenStudy (ybarrap):

Ok, so you do know the formula for the sum of an infinite geometric series, it's actually one easy one to remember once you know it.

OpenStudy (ybarrap):

What is the 1st term?

OpenStudy (anonymous):

im not sure the formula.

OpenStudy (anonymous):

well do you mean the lim n->inf r^n = 1 if r<1 and 0 if |r|<1

OpenStudy (ybarrap):

$$ \Large{ \sum_{n=0}^\infty 2\left ( \cfrac{2}{3}\right )^n\\ =2\times\left ( \cfrac{2}{3}\right )^0+2\times\left ( \cfrac{2}{3}\right )^1+\cdots\\ =2\times1+2\times\left ( \cfrac{2}{3}\right )+\cdots\\ =2+\cfrac{4}{3}+\cdots\\ } $$ Make sense so far?

OpenStudy (ybarrap):

Do you see the 1st term?

OpenStudy (ybarrap):

The terms in parenthesis above is the ratio

OpenStudy (ybarrap):

The formula for the sum uses this ratio: $$ \sum_{n=0}^\infty 2\left ( \cfrac{2}{3}\right )^n\\ =2\left ( \cfrac{1}{1-a}\right )\\ $$ Where \(a\) is the ratio you found above.